OpenStudy (anonymous):

Solve for x: 5(1-x)-2=3x-4 ...With steps please thank you :)

OpenStudy (mattfeury):

I'll help you through the steps, ok? If you get lost, just let me know where. Order of operations time: Distribute the 5 inside of the parenthesis.

OpenStudy (anonymous):

Okay I did Mattfeury

OpenStudy (mattfeury):

what did you get? write it out here so I can go from there

OpenStudy (anonymous):

(5-5x)-2=3x-4

OpenStudy (mattfeury):

excellent! now combine like terms. we can drop the parenthesis because there is nothing we can do inside of it (it is already simplified). so we have 5 - 5x -2 = 3x - 4 we want all the numbers with x's on one side, and all the just numbers on the other.

OpenStudy (mattfeury):

do you see how we would do this?

OpenStudy (anonymous):

hmm...would it be something like this : 5x-3x=5-2-4 ?

OpenStudy (mattfeury):

close. but remember, what you do to one side of the equation must be done to the other as well. so if we want to get rid of the x on the left side, we would have to add 5x to each side: 5 - 5x - 2 + 5x = 3x - 4 + 5x now we can simplify that

OpenStudy (anonymous):

so it would be 5-2=3x + 5x which would equal 7/8x

OpenStudy (anonymous):

?

OpenStudy (mattfeury):

that's right, but it looks like you forgot to write the +4 in there as well, which is how you get '7'. 7 = 8x

OpenStudy (anonymous):

Opps yes I meant to put the 4 ha...Thank you so much for the help :)

OpenStudy (mattfeury):

no problem! just for closures sake, divide by 8! x = 7/8 excellent work! i will definitely give you a medal for trying so hard and not just looking for an answer :)

OpenStudy (anonymous):

:) :) Thanks!